Weighted Average Grade (missing grade). Suppose that your grade for the class is dependent upon four marking periods (MP1, MP2, MP3, MP4), each worth 20%, a midterm (MT) worth 10%, and a final exam (FE) worth 10%. Also, you have yet to take the final exam, and you want to earn a 90 average for the year. Your grades so far are given below. What score do you need to earn on the final exam in order to obtain a 90 for the year?

Grade/Term MP1 89 MP2 93 MT 85 MP3 88 MP4 94 FE
1
Expert's answer
2020-10-21T15:54:16-0400

The weighted average is

"89 \cdot 0.2 + 93 \cdot 0.2 + 85 \cdot 0.1 + 88 \cdot 0.2 + 94 \cdot 0.2 + 0.1 \cdot X =90"

where X is a grade for FE.

"81.3 + 0.1 \cdot X =90"

"0.1 \cdot X =8.7"

"X = 87"

Answer: 87